OpenStudy (anonymous):

Which of the following is the equation of a line parallel to the line y = 4x + 1, passing through the point (5,1)?

OpenStudy (anonymous):

is it 4x+y=-19?

OpenStudy (anonymous):

You must first use the point (5, 1) for x and y, and then 4 for m as the slope as parallel lines will have equivalent slopes. 1 = 4(5) + b Solve for b to get the y-intercept 1 = 20 + b -19 = b

OpenStudy (anonymous):

So y = 4x - 19 is the equation.

OpenStudy (anonymous):

it has to be in standard form

OpenStudy (anonymous):

Rearrange in that case to ax+by+c = 0

OpenStudy (anonymous):

4x - y - 19 = 0

OpenStudy (anonymous):

so, in Ax+By=C, it would be 4x-y=19?

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

thanks a lot! :D

OpenStudy (anonymous):

Np.